If a card is picked at random from a standard 52-card deck, what is the probability of getting a red card or a black face card?
Answer #1

There are 26 red cards and 6 black face cards out of 52. So we have

P(red) = 26/52

P(black face card) = 6/52

Since events "red card" and "black face card" are mutually exclusive so the probability of getting a red card or a black face card is

P(red or black face card) = P(red) + P(black face card) = 26/52 + 6/52 = 32/52 = 8/13 = 0.6154

Answer: 8/13 or 0.6154